OpenStudy (anonymous):

Simplify: A. 13/8 B. 8/13 C. 32/84 D. 64/169

jhonyy9 (jhonyy9):

A. 13/8 B. 8/13 C. 32/84 =16/42 =8/21 D. 64/169

OpenStudy (anonymous):

thanks

jhonyy9 (jhonyy9):

yw good luck bye